### v1

# Robustifying Long-term Human-Robot Collaboration through a Hierarchical and Multimodal Framework

Abstract
Long-term Human-Robot Collaboration (HRC) is crucial for developing flexible manufacturing systems and for integrating companion robots into daily human environments over extended periods. However, sustaining such collaborations requires overcoming challenges such as accurately understanding human intentions, maintaining robustness in noisy and dynamic environments, and adapting to diverse user behaviors. This paper presents a novel multimodal and hierarchical framework to address these challenges, facilitating efficient and robust long-term HRC. In particular, the proposed multimodal framework integrates visual observations with speech commands, which enables intuitive, natural, and flexible interactions between humans and robots. Additionally, our hierarchical approach for human detection and intention prediction significantly enhances the system’s robustness, allowing robots to better understand human behaviors. The proactive understanding enables robots to take timely and appropriate actions based on predicted human intentions. We deploy the proposed multimodal hierarchical framework to the KINOVA GEN3 robot and conduct extensive user studies on real-world long-term HRC experiments. The results demonstrate that our approach effectively improves the system efficiency, flexibility, and adaptability in long-term HRC, showcasing the framework’s potential to significantly improve the way humans and robots work together.
Index Terms:
Human-robot collaboration; Multimodality; Intention prediction; Hierarchical planning; Assembly.
I INTRODUCTION
Human-robot collaboration (HRC) has emerged as a pivotal area of research, particularly as robots become more prevalent in human-centric environments. Although frameworks and technologies supporting short-term HRC have advanced significantly[1, 2], long-term HRC is still not as well developed [3]. In industrial settings, utilizing long-term HRC can significantly enhance productivity and adaptability in complicated manufacturing processes [4]. In-home environments, robots acting as companions or elderly care agents can improve quality of life through long-term HRC [5]. However, achieving effective and robust long-term HRC presents several challenges: 1) Accurate Understanding of Human Intentions: Robots must accurately interpret human actions and intentions in complex and long-term tasks. 2) Robust to Noise: Systems must remain robust against environmental noise and unexpected disturbances. 3) Efficient Collaboration: Interactions should be optimized in completing time for varying contexts. 4) Adaptable to Different Users: Systems should adapt to individual users with diverse behaviors and preferences. .
Figure 1: Overview of the hierarchical task graph. The task graph represents the long-term HRC task, such as toy car assembly (as shown on the left). At the higher level (middle section), each node denotes a subtask. Each subtask consists of a sequence of collaborative actions between the human and the robot (as shown on the right).
Addressing these challenges is essential for the success of long-term HRC in real-world applications. While existing research has made contributions in individual aspects of challenges, there remains a critical need for integrated solutions that address several challenges simultaneously. For instance, studies such as [1, 2, 6] explore adaptable HRC through human motion prediction. However, these works often overlook the robustness to disturbances and the seamless integration required for effective human-robot collaboration. Similarly, to enhance the understanding of human intentions, research by [7, 8] explores multimodal HRC frameworks; however, these frameworks often depend on additional wearable devices. This reliance on wearable devices can limit the generalizability of these systems across different interaction contexts and user populations. Furthermore, some previous works often fail to integrate these techniques into a unified framework and lack real-world demonstrations of complex long-term HRC tasks [9].
In this work, we focus on enhancing long-term HRC, characterized by long sequences of interactions. We utilize a hierarchical task graph to illustrate these long-term HRC tasks. The task graph used in our experiment is depicted in fig. 1. The aim for humans and robots is to progress from the initial node (at the bottom ) to the completion node (at the top), which signifies the end of the task, through a series of interactions. Accurate prediction of human intentions is critical for the robot to determine its collaborative controls effectively. From the task graph perspective, predicting human intentions equates to forecasting the subsequent node in the task graph. Given the long sequences of these interactions, long-term HRC tasks are prone to compounding errors; therefore, predictions must be precise, resilient to disruptions, and tailored to the diverse behaviors of individual humans. This capability ensures that the robot can proficiently plan and execute collaborative controls.
To address these challenges, we have developed a multimodal framework that integrates both visual and auditory signals to enhance the accuracy of intention prediction and current-position estimation within the task graph. The reliance on a single modality often fails to provide accurate predictions over complex, long-term tasks. Further improvements include a hierarchical design in both the human pose detection and intention prediction modules. This design strategy enables the system to initially predict the human’s actions and subsequently refine these predictions using the task graph to correct the detected actions and the corresponding node positions. This hierarchical approach not only increases the precision of intention understanding but also minimizes the impact of environmental noise and disturbances. By combining multimodal observations with hierarchical prediction, our proposed framework aims to achieve robust, efficient long-term HRC.
In summary, our main contributions are as follows:
  1. 1.
    We present an effective framework for solving long-term human-robot collaboration tasks.
  2. 2.
    We introduce a hierarchical pose detection algorithm that effectively mitigates disturbances in multi-human environments, ensuring reliable human detection.
  3. 3.
    We validate our framework in real-world complex toy car assembly tasks. Through physical experiments and user studies, we demonstrate that our approach significantly improves the robustness, time efficiency, and flexibility of human-robot collaboration.
II RELATED WORKS
Long-term Human-Robot Collaboration. Previous research on human-robot collaboration (HRC) has primarily focused on short-term tasks with long-horizon action sequences[10, 11], rather than fully addressing long-term, complex tasks. While these studies explore important aspects of task planning over extended action sequences [12, 13], there is limited work specifically targeting long-term human-robot collaboration that achieves comprehensive, goal-oriented tasks. Pirk et al.[10] successfully recognized long-term human sub-assembly actions but did not demonstrate human-robot collaboration in real-world scenarios. In this work, we implement our framework on a real-world long-term complex assembly task.
Multimodal Human-Robot Collaboration. The interaction methods of the multimodal HRC framework can be summarized into four categories: vision, auditory and/or language, physiological sensing (human-centered sensing), and haptics (robot sensing) [8]. Physiological sensing [14, 8] typically requires extra wearable devices to detect bioelectrical signals, which often reduces generalizability across different interaction contexts and increases the operational complexity. Vision-language fusion-based multimodal human-robot collaboration frameworks have proven effective in tasks such as navigation and social interaction [15, 16]. Liu et al. [7] demonstrated the effectiveness of model fusion of vision and auditory modalities in assembly tasks, although their approach has not been tested on real robots. Maurtua et al. [17] relied on human gestures for vision perception and auditory input to conduct human-robot collaboration in industrial environments. In this work, we fuse vision and auditory modalities, requiring only a camera and a microphone, in assembly tasks with potential applications for industrial and home service robots.
Human Pose Detection and Intention Prediction. High-fidelity and robust human behavior prediction is a key component of achieving safe HRC [18, 19]. Several approaches utilize human detection models, such as OpenPose [20], to identify human key points, which are then used to predict human behaviors including intentions and/or trajectories [2, 6]. Reducing disturbances in human pose detection has been extensively studied using signal filtering techniques [21, 22]. Traditional filtering methods primarily mitigate temporal disturbances [23]. In contrast, we focus on spatial aspects and propose a novel hierarchical framework to reduce disturbances in HRC. Additionally, we implemented an online model adaptation algorithm to adapt the prediction models to accommodate different users’ behaviors [24].
III PROBLEM FORMULATION
The Long-Term HRC Task. We consider a scenario where a human $A_{H}$ collaborates with a robot $A_{R}$, to achieve a set of predefined goals within a long-term task $\mathcal{G}$. The ultimate goal, $g$, is the completion of the task, represented by reaching the top of the hierarchical task graph $\mathcal{G}$, as shown in fig. 1. We define $\mathcal{G}$ as a Hierarchical and Multi-Agent Temporal Plan Graph (TPG) [25], where task sequences are structured within a directional graph. The nodes $\mathcal{N}$ represent individual sub-tasks, and the directed edges denote the precedence of these nodes.
Figure 2: Example of a temporal plan graph for each subtask. Note that, for a given subtask, there may be multiple plans that can achieve the same goal. In this figure, two different human-robot plans (action sequences) are illustrated to achieve the same subtask from the same initial state.
Sub-task Configuration. At the highest level, $\mathcal{G}$ is the overarching long-term HRC task. Within this framework, each node corresponds to shorter-term HRC tasks, referred to here as sub-tasks. Each subtask $N_{i}\in\mathcal{N}$ is also organized as a multi-agent TPG, which represents both the precedence and collaborative dynamics between human and robot actions [26]. Each vertex at the lowest hierarchy level represents an action performed by a human or robot. Figure 1 illustrates a three-level structure of the task graph—from the multi-agent long-term task level to the multi-agent short-term sub-task level, down to the single-agent shortest-term action level. A further depiction of the lowest level temporal plan graph is shown in fig. 2, emphasizing that in HRC, the robot must predict human intentions, which involves foreseeing the subsequent node a human might conduct based on their behavior’s inherent property.
The Objective Function. Let $E$ denote the environment in the HRC task, incorporating both humans and objects. Let $P$ denote the task progress, which tracks the completion of subtasks (or nodes within the task graph), and updates as the task advances. At timestep $t$, the human’s intention $x_{H}^{t}$ is influenced by the current state of the environment $E^{t}$ and task progress $P^{t}$. The robot predicts human intentions, $\hat{x}_{H}^{t}$, using its learnable model $\hat{f}_{H}$:
$\hat{x}_{H}^{t}=\hat{f}_{H}(E^{t},P^{t}),$ (1)
The predicted intentions guide the robot in generating a motion sequence $q=[q^{1},\cdots,q^{n}]$. The objective for the robot is to find an efficient and robust motion sequence $q$ under safety and dynamics constraints. Let $J_{g}$ represent the task-specific cost of achieving the goal $g$, such as a tracking cost. To reflect time efficiency, we include the task completion time $T_{g}$, weighted by $\lambda$, in the objective function. Thus, the robot’s objective in HRC is formulated as:
$\min_{q=[q^{1},\cdots,q^{n}]}~{}J_{G}(q,g)+\lambda T_{g}(q,g)$ (2)
[…]
where $M_{R}(\cdot)$ represents the robot motion planning function. Accurate and robust prediction of human intentions, $\hat{x}_{H}^{t}$, is therefore pivotal to the process.
IV PROPOSED FRAMEWORK
Figure 3: The architecture of the proposed HRC framework.
To address these challenges of long-term HRC tasks as discussed in section I, we propose a hierarchical and multimodal HRC framework, as illustrated in fig. 3. The framework includes two perception modules from different modalities (a visual perception module and an audio perception module), four modules for task planning (pose detection, intention prediction, speech recognition, and robot motion planner), and a robot motion controller.
  • Perception. Visual Perception Module captures visual information from the environment and humans using RGBD cameras. Audio Perception Module processes audio information through microphones.
  • Planning. Pose Detection Module is responsible for identifying human poses and extracting key point trajectories. Its hierarchical design allows the framework to mitigate disturbances in complex, multi-human environments, ensuring more robust performance. Intention Prediction Module predicts human intentions based on the trajectories of human keypoints and task progress. Similarly, the hierarchical design of this module enables robots to respond to human actions more robustly. Speech Recognition Module recognizes speech commands for audio-based robot motion generation and corrects errors in robot motion caused by inaccuracies in human intention prediction. Robot Motion Planner Module outputs the high-level robot motions based on the fusion of predicted human intention, recognized human speech commands, and task progress. Figure 4 presents a detailed overview of the framework’s hierarchical planning modules.
  • Control. Controller Module receives the high-level motions from the planner module and translates them into low-level robot actions for execution.
Figure 4: The overview of the planning modules.
IV-A Multimodal Perception
In human-to-human interactions, visual observations alone often fail to fully capture human intentions, necessitating the integration of verbal communication for comprehensive context. Applying this principle to long-term HRC tasks, we propose a multimodal framework, as depicted in fig. 3. This framework leverages visual observation to detect human poses and predict human intentions. However, visual data alone can be ambiguous, as the same human pose may indicate different intentions depending on the context. To resolve these ambiguities, we incorporate audio observations to add more contextual information. Speech commands enable humans to make clarifications and adjustments to the robot’s motions, informed by its visual interpretations. This integration yields a more comprehensive understanding of human intentions and enhances the robot’s adaptability to dynamic user demands. The multimodal approach offers two significant advantages: 1) By integrating visual and audio signals, the robot achieves a comprehensive perception of human intentions, allowing for more complex and long-term collaborations. 2) The inclusion of speech enables more natural and flexible interactions, making the system more user-friendly and accessible.
We employ mutual information analysis to demonstrate the effectiveness of our multimodal framework. Let $O_{V}$ denote the robot’s visual observation from the environment $E$, captured by the RGBD camera; and $O_{A}$ denote the robot’s audio observation, captured by the microphone. The robot’s performance in achieving the goal $g$ is enhanced by these informative observations. We consider the mutual information between the goal and observations:
$I(g;O_{V},O_{A})=H(g)-H(g|O_{V},O_{A})$ (4)
where $I(\cdot,\cdot)$ denotes mutual information and $H(\cdot)$ denotes entropy. By the nature of information fusion, we have:
$H(g|O_{V},O_{A})\leq\min\{H(g|O_{V}),H(g|O_{A})\}$ (5)
This inequality holds because the combined information from fused modalities $(O_{V},O_{A})$ is no less than the information from single modality $O_{V}$ or $O_{A}$ individually. We also provide the proof in section -A. Combining the above expressions, we find:
$I(g;O_{V},O_{A})$ $\geq\max\{H(g)-H(g|O_{V}),H(g)-H(g|O_{A})\}$
$=\max\{I(g;O_{V}),I(g;O_{A})\}$ (6)
Therefore, we conclude that the multimodal framework yields greater mutual information between the goal and the observations than a single-modal framework, significantly enhancing the robot’s capability to handle complex and long-term collaboration tasks.
IV-B Hierarchical Pose Detection
Effective long-term HRC critically depends on accurately predicting human intentions, which are deduced from the estimated human poses obtained from visual observations. However, visual observations can often be noisy, leading to errors in pose detection and subsequent intention prediction, ultimately impacting the robot’s motion control. This noise can arise from various sources, such as the presence of multiple humans in the scene or cluttered backgrounds. To mitigate these disturbances, we propose a hierarchical pose detection module that enhances both accuracy and efficiency. Our module operates through a two-step approach, based on the distance between the human and the robot, as outlined in the grey box of fig. 4. The first step involves identifying all humans in the scene using real-time RGB images from visual observations. Concurrently, a depth map, from the depth camera, estimates the distance of each detected human from the robot. If no human is detected within a predefined close range (e.g., 2 meters), the detection cycle repeats, and downstream HRC tasks are temporarily paused. This step ensures that the robot focuses on humans who are likely to interact closely, based on their proximity. Once a human is detected within the effective range, the RGB image is cropped to focus on the nearest individual. This cropped image is then processed by the pose estimation model to detect specific poses and key points. A sample outcome of this process is displayed in fig. 7. This hierarchical structure offers two main advantages: 1) By activating the pose estimation and planning modules only when a human is detected within a practical range, the system conserves computational and energy resources, operating more efficiently. 2) The hierarchical filtering effectively reduces the interference from non-target humans and background noise, thus improving the precision of pose detection. The following equation represents the operation of our proposed hierarchical pose detection model $\hat{f}_{\rm detection}$, which estimates the human pose $\hat{z}_{H}^{t}$ from visual observations: $\hat{z}_{H}^{t}=\hat{f}_{\rm detection}(O_{V}^{t})$.
IV-C Hierarchical Human Intention Prediction
Hierarchical human intention prediction is fundamental to effective long-term human-robot collaboration (HRC). This process involves two interconnected levels of intention analysis: human action prediction and human plan prediction. Action prediction focuses on low-level intentions that dictate short-term motion patterns, such as ”fetching a screw.” At a higher level, plan prediction addresses long-term intentions that are structured as sequences of such actions. As shown in the green box of fig. 4, the robot first predicts the human action from the current human pose. Then the high-level intention $\hat{x}_{H}^{t}$, or the human plan, considers both the current task progress and task graph.
IV-C1 Action Prediction
Figure 5: Action prediction pipeline
Figure 5 illustrates our action prediction pipeline. Our approach does not simply predict human actions from the current pose trajectory. Instead, we initially predict the future trajectory and subsequently infer the action based on both the current and predicted trajectories. This method ensures that the action prediction is informed by anticipated future movements. Research has demonstrated that joint training of the action and trajectory prediction models enhances performance compared to training standalone action prediction models [2]. We train both the trajectory and action prediction models using supervised learning, utilizing data collected offline. The trajectory prediction model is optimized using regression loss (specifically L2 loss for predicted trajectories), while the action prediction model employs classification loss (namely cross-entropy loss for predicted actions). Once trained, the action prediction model $\hat{f}_{\text{action}}$ applies to predict human actions $\hat{x}_{H,\text{action}}^{t}$ based on the current human pose: $\hat{x}_{H,\text{action}}^{t}=\hat{f}_{\text{action}}(\hat{z}_{H}^{t}).$
In long-term HRC, prediction models designed for one human often fail to generalize across different users due to variations in personal preferences and behavioral styles. To address this, we implement an online adaptation method during deployment [18]. This approach allows the robot to continually refine its predictions by incorporating real-time behavior observations of each user [1]. Our joint training framework facilitates the dynamic adaptation of the action prediction model to various users in real-time HRC scenarios. Specifically, the trajectory prediction model is optimized online using observed trajectory feedback. Since the action prediction depends on the trajectory prediction module, any updates to the trajectory model also enhance action prediction accuracy [22]. For more details on this online adaptation approach in this paper, refer to the work [24]. This technique significantly improves prediction performance, ensuring that the robot’s responses accurately align with each user’s behaviors and intentions.
IV-C2 Plan Prediction
After obtaining predicted action $\hat{x}_{H,\text{action}}^{t}$, the robot predicts the high-level intention $\hat{x}_{H}^{t}$, or the human plan. his involves integrating the current task progress $P_{t}$, which encapsulates the sequence of robot actions executed thus far, with the task graph $\mathcal{G}$. The robot identifies the reference action sequence $R^{\ast}$ within the task graph that most closely aligns with the recorded task progress. This alignment is formulated as:
$R^{\ast}=\arg\min_{R\in\mathcal{G}}d(P^{t},R),$ (7)
where $d$ measures the distance between the observed task progress $P_{t}$ and each possible reference sequence $R$ within the task graph. The distance metric $d$ is calculated using the dynamic time warping (DTW) algorithm, renowned for its capability to flexibly align sequences with varying lengths and timings [27]. The result, $R^{\ast}$, represents the best-matched reference action sequence. Note that $R^{\ast}$ may not be unique, since the multiple action sequences may present the same distances due to the potential for the current task progress to align with different action sequences leading to the goal. Following this alignment, the human plan prediction is obtained by refining the predicted human action based on the most probable action class within $R^{\ast}$. This adjustment is expressed as:
$\hat{x}_{H}=R^{\ast}(\hat{x}_{H,\text{action}}^{t}),$ (8)
where $\hat{x}_{H}$ represents the predicted human plan. This equation aligns $\hat{x}_{H,\text{action}}^{t}$ with the most likely action within the set of potential action sequences in $R^{\ast}$. This hierarchical framework effectively integrates low-level human action predictions with high-level task plans. By ensuring that the robot’s predictions are consistently aligned with the task graph and progress, it significantly enhances the accuracy and reliability of human-robot collaboration.
IV-D Speech Recognition and Robot Motion Planning
For the audio modality $O_{A}$, the robot recognizes the audio-based human intention or speech command, $\hat{s}_{H}^{t}$, using a pretrained speech recognition model $\hat{f}_{\rm speech}$: $\hat{s}_{H}^{t}=\hat{f}_{\rm speech}(O_{A})$. In deployment, this model’s parameters are fixed. To enhance the quality of speech recognition, a Voice Activity Detector (VAD) is utilized to filter the input. This VAD activates the speech recognition module only when significant audio signal energy, not noise, is detected.
Upon acquiring the vision-based human intention $\hat{x}_{H}^{t}$ and the speech command $\hat{s}_{H}^{t}$, the robot’s motion planner generates the robot motions from both modalities as follows:
$q_{V}^{t:t+m}$ $=M_{R}(\hat{x}_{H}^{t},P^{t}),~{}~{}q_{A}^{t:t+m}=M_{R}(\hat{s}_{H}^{t},P^{t}),$ (9)
$q^{t:t+m}$ $=(1-w^{t})q_{V}^{t:t+m}+w^{t}q_{A}^{t:t+m},$ (10)
where $q_{V}^{t:t+m}$ and $q_{A}^{t:t+m}$ represent the motion sequences planned based on vision-derived intentions and audio commands, respectively. The final robot motion, $q^{t:t+m}$, integrates these motions, weighted by a scalar $w_{t}$. In our work, we implement an indicator function that gives priority to audio-based observations. Specifically, if keywords are recognized in the speech recognition output, we set $w^{t}=1$, engaging the audio-based motion planner exclusively. Otherwise, the vision-based motion planner predominates, and $w^{t}$ is set to zero. The motion controller operates as a feedback controller with an update frequency of approximately 30 Hz.
V EXPERIMENTS
V-A Toy Car Assembly Task
We evaluate the effectiveness of our proposed framework through a long-term HRC task, specifically the toy car assembly task. This assembly process incorporates various components, including both short and long tubes, connectors, screws, wheels, and panels. This task is divided into four stages: assembling the bottom, the middle, the top of the car, and installing the wheels, as shown in fig. 1. For simplicity, we will refer to the first three stages as The Bottom, The Middle, and The Top. The assembly order for The Middle and The Top is flexible, allowing variations in the sequence order. Each stage is designed to require collaborative efforts between the human and the robot. The actions and motion sequences within these stages are non-deterministic, offering a range of valid action sequences that can successfully complete each sub-task. This flexibility means that the specific sequence of actions may vary according to individual human preferences. In this scenario, the robot’s role is to support the human collaborator by delivering necessary objects and performing specific assistive actions, such as rotating or lifting objects.
As shown in the bottom part of fig. 3, the tools and objects needed for the assembly task are positioned across different locations. Connectors, screws, and other small materials are placed on a table near the human to allow easy access. Conversely, the tubes are stored on a table that is closer to the robot arm and relatively farther from the human. During the assembly process, the robot assists based on the human’s intentions and the current task progress. For instance, if the human intends to reach for connectors, the robot responds by delivering the appropriate tube—either short or long—based on the observed signals. When the human intends to retrieve screws, the robot aids by spinning the base-object to facilitate easier handling. Similarly, when the task involves attaching wheels, the robot assists by lifting the car’s body to better position it for wheel installation. The source code and demonstration video for the HRC toy car assembly task are available at https://github.com/intelligent-control-lab/Robust-Hierarchial-Multimodal-HRC.
V-B Experimental Design
V-B1 Implementation Details
We test our Human-Robot Collaboration (HRC) framework using the KINOVA GEN3 robotic arm, equipped with an OAK-D Lite camera for video (capturing both RGB and depth maps) and a directional microphone for audio. Below, we detail the neural network models used across various components of our framework, noting that only the action prediction network is trained; all others utilize pretrained weights.
  • Human (Object) Detection. We use a pretrained MobileNet model [28] to detect and classify 21 object types from RGB data. For our purposes, we specifically focus on ”human” detections, applying a confidence threshold of 0.6.
  • Human Pose Detection. The pretrained BlazePose model [29] extracts 33 key points, of which we focus on 15 upper-body points for action and intention prediction.
  • Human Action Prediction. We implement the DLinear model [30] for action and trajectory prediction, which decomposes input time series into trend and seasonal components. This model is trained using offline data.
  • Speech Recognition. The DeepSpeech model [31], an open-source Speech-to-Text engine is utilized for recognizing and processing speech commands.
V-B2 Hypothesis
We evaluate the effectiveness of the proposed multimodal hierarchical HRC framework by testing five hypotheses:
  • H1: The framework enhances the efficiency of HRC.
  • H2: The framework reduces disturbances in multi-human environments and increases action prediction accuracy.
  • H3: The framework increases task completion rates in collaborative tasks.
  • H4: The framework adapts to varying human behaviors.
  • H5: The framework improves user flexibility and satisfaction in collaboration tasks.
V-B3 User Study
We conducted a user study with 10 participants, focusing on 3 tasks from a toy car assembly problem. The study included three experimental settings: vision-only, audio-only, and multimodal, each repeated 4 times per participant. The tasks involved two assembly stages and two robot actions:
  • Task 1: Pick and place objects based on human intentions to complete The Bottom.
  • Task 2: Rotate the car’s bottom to assist the human in screwing in The Bottom.
  • Task 3: Collaborate with the human using the hierarchical framework to complete The Middle.
V-B4 Dataset
We collected an offline dataset to train and evaluate the action prediction module. The training data includes 3,003 trajectories from two users, while the test data consists of 2,088 trajectories from these two users and two additional users to evaluate robustness. Each sequence of trajectories contains 5 time steps. The five hypotheses are evaluated using offline test data, user studies, or both.
V-C Results
V-C1 H1 - Time Efficiency
Table I compares the average completion times of three different frameworks—each utilizing different modalities—across the three sub-tasks outlined in the user study section V-B3. The table summarizes results from 40 trials for each task and modality. The percentage values in the table are calculated relative to the longest completion time recorded for each task. It is important to note that in Task 3, human intention cannot be accurately predicted using vision alone, as human trajectories exhibit minimal movement during the screwing process. This limitation prevents the intention prediction model from making accurate predictions based solely on visual input, resulting in an inability to complete Task 3 with the vision-only framework. Thus, an infinite completion time is assigned for this scenario. As the results show, the multimodal framework achieves the shortest completion time across all tasks compared to the single-modal frameworks. This demonstrates that the proposed framework effectively enhances the efficiency of human-robot collaboration.
Vision Audio Multimodal
Task1 (%) $75.70\pm 10.56$ $72.02\pm 9.10$ $\textbf{67.42}\pm\textbf{4.91}$
Task2 (%) $48.42\pm 20.90$ $43.72\pm 13.50$ $\textbf{39.45}\pm\textbf{14.83}$
Task3 (%) $\infty$ $74.19\pm 10.83$ $\textbf{65.83}\pm\textbf{3.86}$
TABLE I: The average completion time with standard deviation.
V-C2 H2 - Disturbance Minimization
The presence of multiple humans in the same environment can introduce errors in human detection. As discussed in section IV-B, we propose a hierarchical human detection module to address the challenges and disturbances caused by multiple humans. For comparison, sample results from a naive human pose detection model (without hierarchical detection) are shown in fig. 6. As illustrated, the naive detection model can result in errors or misdetections when multiple humans are present. In contrast, the proposed hierarchical framework reduces these disturbances by prioritizing the human closest to and interacting with the robot, as demonstrated in fig. 7.
Figure 6: Misdetection in naive human pose detection model.
Figure 7: Correct detection using the hierarchical human detection framework.
Hierarchical Human Detection. We conducted the following case study to numerically evaluate the effectiveness of the hierarchical framework. In this setup, one person stood close to the robot, serving as the human operator, while another person walked through the environment to introduce noise for the detection model. To evaluate the framework’s robustness, we measured the keypoint deviation over time. Figure 8 shows the keypoint deviation for three methods: naive detection without hierarchical human detection, naive detection with Kalman filtering, and hierarchical human detection. The ground truth is represented by a constant zero line (green dotted line). As illustrated, the naive detection method, even when smoothed with Kalman filtering, results in sharp fluctuations in the detected keypoints, with a variance of 0.013. In contrast, the proposed hierarchical human detection method (black curve) provides robust and stable detection, closely aligning with the ground truth. These results support the hypothesis that the hierarchical framework effectively minimizes disturbances in multi-human environments.
Figure 8: Keypoint deviation with and without hierarchical human detection.
Action prediction. Table VI compares the action prediction accuracy of the proposed hierarchical human detection framework against a naive baseline and four baselines using filtering methods. The methods are: (a) ”w/o hierar,” which predicts actions using a naive human pose; (b) ”median/wiener/kalman/ema filter,” which applies respective filters to the naive human pose before predicting actions; and (c) ”w/ hierar,” which uses the hierarchical detection framework for action prediction. In the table, ”Action1” corresponds to ”get_connectors,” ”Action2” to ”get_screws,” and ”Action3” to ”get_wheels.” ”No Action” indicates instances where the human’s action did not match any of these three actions. ”All” represents the overall average accuracy across all actions. The results clearly show that hierarchical human detection significantly improves action prediction accuracy compared to other methods.
w/o hierar median filter wiener filter kalman filter ema filter w/ hierar
No Action 0.5824 0.4291 0.4713 0.2835 0.3295 0.9127
Action1 0.9810 0.8659 0.8644 0.5729 0.7930 0.9985
Action2 0.2196 0.6738 0.7234 0.8095 0.6211 0.9711
Action3 0.3835 0.6627 0.6406 0.7992 0.8996 0.9200
All 0.5139 0.7001 0.7190 0.6811 0.6971 0.9617
TABLE II: Comparison of action prediction accuracy.
To minimize critical class-dependent misclassification errors—such as mistakenly classifying ”stay constant” as ”reach out to human”—we implement sensitivity-aware intention prediction. This approach strategically ”transfers” high-sensitivity errors to lower-sensitivity ones, improving overall prediction accuracy. Further details are available in section -C.
V-C3 H3 - Task Completion
As discussed in section IV-A, a multimodal observation framework significantly enhances the information captured from the environment. This approach reduces errors compared to single-modality perception systems, which often lack adequate information about humans and the environment. We evaluated the effectiveness of multimodal observations during the user study. Table III shows the task completion rates for three tasks in the user study, conducted under three different observation modalities. Note that Task 3 cannot be completed using a vision-only framework, as explained in section V-C1. As shown intable III, the multimodal framework allows the robot to respond more proactively by utilizing the richer information it gathers. This enables the robot to complete tasks more effectively and better meet user expectations.
Vision Audio Multimodal
Task1 (%) $70.63\pm 18.65$ $90.63\pm{6.75}$ $\textbf{93.75}\pm\textbf{9.32}$
Task2 (%) $68.75\pm 22.24$ $63.75\pm 18.35$ $\textbf{85.38}\pm\textbf{13.58}$
Task3 (%) 00 $81.25\pm 19.98$ $\textbf{96.25}\pm\textbf{4.37}$
TABLE III: Task completion rates under different modalities.
V-C4 H4 - Adaptibility
We employ an online adaptation strategy for the trajectory prediction model to refine its predictions. We conducted experiments to evaluate the performance of an offline-trained prediction model across four different users. Accuracy was used to evaluate intention prediction, while mean-squared-error (MSE) was used to evaluate trajectory prediction. Table IV presents the results of predictions with and without online adaptation. The results clearly show that online adaptation improves the performance of both intention and pose prediction. The improvement in trajectory prediction is particularly significant, with a 24% reduction in MSE. This is expected, as different individuals may have varying pose and pose preferences. The results demonstrate that the proposed framework effectively adapts to different human behaviors.
w/o adaptation with adaptation
Intention Accuracy 0.9617 0.9626
Pose MSE 1.2574 1.017
TABLE IV: Prediction results with and without online adaptation.
V-C5 H5 - User Satisfaction
We conducted a user study to evaluate user satisfaction with the toy car assembly HRC task, collecting responses to evaluate the effectiveness of the proposed multimodal framework. Figure 9 participants’ responses to six questions across three different HRC frameworks: vision-based, audio-based, and the proposed multimodal framework. As shown, the percentage of participants who selected ”strongly agree” was significantly higher for the multimodal framework compared to the single-modal frameworks. For example in Figure 9(f), many users ”strongly agree” with the statement, ”I am satisfied with this framework,” when using the multimodal framework. Participants rated their agreement on a scale from 1 (strongly disagree) to 5 (strongly agree). The average satisfaction score for each framework across the six questions is summarized in section -B.
Figure 9: User Study
In the user study, the multimodal framework received the highest satisfaction ratings from most participants, outperforming the single-modal frameworks. Specifically, user feedback highlighted that: 1) Users found it easier to interact with the robot using the proposed multimodal framework, indicating an improved user experience. 2) Users expressed a greater willingness to collaborate with the multimodal framework over an extended period, suggesting that it effectively reduces user workload. 3) Overall, users reported higher satisfaction when working with the multimodal framework. This user study demonstrates the effectiveness of the multimodal framework in creating a user-friendly HRC system for long-term HRC tasks.
VI CONCLUSIONS
In this work, we propose a hierarchical and multimodal framework for human-robot collaboration tasks. Vision observations from the environment are used to predict human intentions through hierarchical human pose detection and intention prediction, while audio observations are employed to recognize speech commands. The predicted intentions and recognized speech are then used to guide the robot’s motions according to the plan graph. Experiments and user studies demonstrate that the proposed framework significantly improves the efficiency of long-term HRC tasks and increases user satisfaction.